# Causality-Aware Temporal Projection for Video Understanding in Video-LLMs

### v2

@@ II-A Video-Language Models
Video-Language Models (Video-LLMs) extend the instruction-following capabilities of LLMs [1, 21] to the video domain by coupling a frozen vision encoder with a lightweight projection module. Early systems such as Video-ChatGPT [20] and Video-LLaMA [28] established this paradigm using linear or shallow projectors to map visual features into the LLM token space. Subsequent works, including Video-LLaVA [16], further improved multimodal alignment through instruction tuning and better visual-language pretraining. More recent models such as LLaVA-NeXT [29] and Qwen2-VL [22] scale video understanding by increasing visual granularity through dynamic resolution and advanced token compression strategies. While these approaches achieve strong performance, temporal modeling is primarily handled implicitly within the LLM’s self-attention layers. As a result, the projection stage is often agnostic to temporal directionality, treating frame-level features as unordered or symmetrically related tokens.
Complementary to video-centric modeling, fine-grained multimodal understanding of complex visual content has been studied in prior work [5, 23], highlighting the importance of structured visual representations for downstream reasoning. Beyond accuracy, recent studies have also examined reliability aspects of multimodal models, including robustness under visual adversarial perturbations [4] and stability under task-incremental learning settings where forgetting can occur [15]. In contrast, our work focuses on parameter-efficient video QA and introduces explicit architectural constraints at the projection stage to preserve temporal ordering, rather than relying on implicit temporal reasoning to emerge within the LLM.
@@ II-B Spatio-Temporal Modeling and Token Compression
Modeling temporal structure is central to video understanding, as it captures motion, event ordering, and causal dependencies. Vision Transformers such as TimeSformer [3] and ViViT [2] explicitly model spatio-temporal interactions, but their computational cost limits their direct adoption as connectors for large language models. Structured attention and relation modeling [10, 27] have been explored in vision transformers to improve efficiency and representation quality. To maintain efficiency, many Video-LLMs compress frame-level features into a compact set of tokens before projection. Recent methods such as ST-LLM [18] and LLaMA-VID [14] reduce redundancy through token pruning or aggregation, but typically rely on global or bidirectional attention mechanisms within the temporal projector. As highlighted by recent benchmarks such as Video-MME [7], the absence of explicit directional constraints can hinder long-range temporal reasoning, particularly in tasks requiring event ordering. Unlike prior approaches that primarily focus on reducing token count, our method jointly addresses what information to preserve and how it is temporally structured. Specifically, Learnable Spatial Aggregation selects salient spatial tokens per frame, while the Causality-Aware Temporal Projector enforces a unidirectional temporal flow consistent with the arrow of time, culminating in a terminal causal sink for global aggregation.

### v3

# V-CORE: Temporally Consistent Video Understanding for Video-LLM

@@ II-A Video-Language Models
Video-Language Models (Video-LLMs) extend the instruction-following capabilities of LLMs [1, 23, 27] to the video domain by coupling a frozen vision encoder with a lightweight projection module. Early systems such as Video-ChatGPT [22] and Video-LLaMA [31] established this paradigm using linear or shallow projectors to map visual features into the LLM token space. Subsequent works, including Video-LLaVA [18], further improved multimodal alignment through instruction tuning and better visual-language pretraining. More recent models such as LLaVA-NeXT [32] and Qwen2-VL [24] scale video understanding by increasing visual granularity through dynamic resolution and advanced token compression strategies. While these approaches achieve strong performance, temporal modeling is primarily handled implicitly within the LLM’s self-attention layers. As a result, the projection stage is often agnostic to temporal directionality, treating frame-level features as unordered or symmetrically related tokens.
Complementary to video-centric modeling, fine-grained multimodal understanding of complex visual content has been studied in prior work [5, 25], highlighting the importance of structured visual representations for downstream reasoning. Beyond accuracy, recent studies have also examined reliability aspects of multimodal models, including robustness under visual adversarial perturbations [4, 17] and stability under task-incremental learning settings where forgetting can occur [16]. In contrast, our work focuses on parameter-efficient video QA and introduces explicit architectural constraints at the projection stage to preserve temporal ordering, rather than relying on implicit temporal reasoning to emerge within the LLM.
@@ II-B Spatio-Temporal Modeling and Token Compression
Modeling temporal structure is central to video understanding, as it captures motion, event ordering, and causal dependencies. Vision Transformers such as TimeSformer [3] and ViViT [2] explicitly model spatio-temporal interactions, but their computational cost limits their direct adoption as connectors for large language models. Structured attention and relation modeling [10, 30, 11] have been explored in vision transformers to improve efficiency and representation quality. To maintain efficiency, many Video-LLMs compress frame-level features into a compact set of tokens before projection. Recent methods such as ST-LLM [20] and LLaMA-VID [15] reduce redundancy through token pruning or aggregation, but typically rely on global or bidirectional attention mechanisms within the temporal projector. As highlighted by recent benchmarks such as Video-MME [7], the absence of explicit directional constraints can hinder long-range temporal reasoning, particularly in tasks requiring event ordering. Unlike prior approaches that primarily focus on reducing token count, our method jointly addresses what information to preserve and how it is temporally structured. Specifically, Learnable Spatial Aggregation selects salient spatial tokens per frame, while the Causality-Aware Temporal Projector enforces a unidirectional temporal flow consistent with the arrow of time, culminating in a terminal causal sink for global aggregation.